OpenStudy (anonymous):

Can someone explain to me why e^ln 7 is equal to 7?

OpenStudy (anonymous):

e and ln are inverses of each other

OpenStudy (anonymous):

Think of an analogy: why does 7 divided by 9 and multiplied by 9 equal 7? Similarly, ln 7 means "e raised to what power is 7?". Then raise e to that power, and you get 7.

OpenStudy (anonymous):

Why isn't it 1 to the seventh power?

OpenStudy (anonymous):

"ln 7" is the value that, when you raise e to it, gives you 7. "e to what power is 7?" is ln 7.

OpenStudy (mertsj):

Because a log is an exponent. So it must be placed on a base. when you write ln 7 you are really saying tell me the exponent that I must put on the base e so that I will get 7. So then you place it on the base e. And WALLA!!! you get 7

OpenStudy (anonymous):

*voila

OpenStudy (anonymous):

I still don't understand.

OpenStudy (anonymous):

Let's use a different base, like 2. What's log (base 2) of 16?

OpenStudy (anonymous):

In other words, 2 to what power equals 16?

myininaya (myininaya):

\[\text{ let } y=e^{\ln(7)}\] we could take natural log of both sides \[\ln(y)=\ln(e^{\ln(7)})\] using properties of log we can rewrite this \[\ln(y)=\ln(7) \cdot \ln(e)\] but ln(e)=1 so we have \[\ln(y)=\ln(7)\] but this would imply y=7

OpenStudy (anonymous):

I get that but i still don't understand the e ln thing

OpenStudy (anonymous):

Do you see why 2 ^ (log base 2 of 16) equals 16?

OpenStudy (anonymous):

@ktklown, yes @myininaya, what is 'e'? and I don't understand how the equation was rewritten. Wouldn't it be ln^2(7) times ln(e)?

OpenStudy (mertsj):

e is a number that is used a lot in calculus and some other classes. It is about 2.7 So when you write ln what you are really writing is log(base 2.7)

OpenStudy (anonymous):

Confucious -- ok, so if you see why 2 ^ (log base 2 of 16) is equal to 16, then, you can probably also see why 2.718 ^ (log base 2.718 of 16) is equal to 16. "e" is just a special number, about 2.71828, and "ln" just means "log base e".

OpenStudy (mertsj):

\[\ln 7 =\log_{e}7 \]

OpenStudy (anonymous):

When I put log(base 2.7) 7 in the calculator I don't get 7. I don't get that when I put in log(base e) 7 either.

myininaya (myininaya):

you shouldn't get 7 if you put that

OpenStudy (anonymous):

How did yo rewrite the equation?

OpenStudy (anonymous):

No, but 2.7 to the POWER of log (base 2.7) of 7 is equal to 7.

myininaya (myininaya):

you have \[e^{\ln(7)}=e^{\log_e(7)}=7\]

OpenStudy (anonymous):

log (base 2.7) of 7 is not 7, it's the number that, if you raise 2.7 to that power, will give you 7.